# ”SNe Ia twins” in the Hubble flow, and the determination of H0

### v1

@@ 5 Methodology
All spectra have been corrected from Galactic reddening. This is done with the purpose of allowing to compare Galactic dust-free spectra. It also helps to learn about dust extinction in the host. The spectra are compared in the frame of the Hubble flow supernova. For that purpose nearby spectra are shifted in wavelength and their flux is diluted in accordance with the (1 + ) law. The shift brings nearby SN Ia to the redshift of the Hubble flow supernova taking into account the heliocentric redshift of nearby SNIa (shifting by the difference, though the redshift of nearby SNIa is always low).
@@
To find the best values and the uncertainties of all variables, we explore the parameter space with Markov Chain Monte Carlo (MCMC) techniques after converting the into a log likelihood function:
where obs corresponds to a given supernova and ref to its reference twin, which represents a whole class. and correspond to the fluxes of the two SNe at different wavelengths, and includes the quadratic addition of the uncertainty on the fluxes of both SNe.
The MCMC allows to get the maximum likelihoood values of each parameters from the posterior distribution of samples, together with their uncertainties, including the propagation of all parameter uncertainties on each parameter distribution.
As we are using different phases, we have a total likelihood function. For two phases, corresponds to 1 and 2 below. Thus the total likelihood function is written as
@@
One can visualize the output of two-dimensional and one-dimensional posterior probability distributions in a corner plot corresponding to 1, 2, 3.
@@ 7.1 The test of the three rung method
We include in Table 15 the distance moduli using the ”SNe Ia twins for life” method as well as those provided by Pantheon+ with SH0ES and by the Carnegie Supernova Project. Those last two methods are calibrated with the Cepheid distances. The SNe Ia moduli in the Hubble flow from the Pantheon+ and SH0ES are different than those used from the CCHP and from those from ”SNe Ia twins”. But they are quite close, except for a disagreement in three SNe Ia. See as well Figure 13).
@@
We have modeled the errors in the distance moduli for our method adding the various contributionss. We have the error in the fit of the distance modulus obtained with the MCMC and the covariance between the modulus and the (E(B-V)), which is very low for our supernovae. We have as well the error due to a difference of stretch between the Hubble flow SNeIa and their nearby twins that goes into the distance modulus. We model it taking into account the CSP impact of the error in stretch in the distance moduli of the SNe Ia. We must say that, in our method, twins at low z and in the Hubble flow have very close stretches: within 0.02 in sBV. The light curves agree so well between low z SNe Ia and their twins in the Hubble flow (once one makes the cosmological (1+z) time broadening and magnitude dilution) that, in their plot of one over the other, they are often indistinguishable (see Appendix).
@@
To calculate the value of H0 we use the expression:
@@
The Pantheon+ data come with their errors in distance moduli in their release and the CCHP distance moduli and errors are generated as in Uddin et al. (2024) code given the entries in the light curve parameterization. We see similar distances in Table 15 by our method when compared with those from Pantheon+ and with Cepheids by the CCHP. There are some differences (in 3 SNe Ia) due to the understimate of reddening, as we will discuss in the next subsection.

### v2

@@ 5 Methodology
All the spectra have been corrected from Galactic reddening. This is done with the purpose of allowing to compare Galactic dust-free spectra. It also helps to learn about dust extinction in the host. The spectra are compared in the frame of the Hubble flow supernova. For that purpose nearby spectra are shifted in wavelength and their flux is diluted in accordance with the (1 + ) law. The shift brings the nearby SN Ia to the redshift of the Hubble flow supernova taking into account the heliocentric redshift of the nearby SNIa (shifting by the difference, though the redshift of nearby SNe Ia is always low).
@@
To find the best values and the uncertainties of all variables, we explore the parameter space with Markov Chain Monte Carlo (MCMC) techniques after converting the into a log likelihood function (the probability of a dataset given the model parameters) that we aim to optimize:
where obs corresponds to a given supernova and ref to its reference twin, which represents a whole class. and correspond to the fluxes of the two SNe at different wavelengths, and includes the quadratic addition of the uncertainty on the fluxes of both SNe. The MCMC allows to get the maximum likelihood values of each parameter from the posterior distribution of samples, together with their uncertainties, including the propagation of all the parameter uncertainties on each parameter.
In fact, as we are using different phases, we have a total likelihood function. For two phases, corresponds to 1 and 2 below. Thus the total likelihood function is written as
@@
One can visualize the output of two-dimensional and one-dimensional posterior probability distributions in a corner plot corresponding to 1, 2, 3. This corner plot shows one and two dimensional projections of the posterior probability distributions of our parameters.
@@ 7.1 The test of the three rung method
We include in Table 15 the distance moduli using the ”SNe Ia twins for life” method as well as those provided by Pantheon+ with SH0ES and by the Carnegie Supernova Project. Those last two methods are calibrated with the Cepheid distances. The SNe Ia moduli in the Hubble flow from the Pantheon+ and SH0ES are quite close to those from the CCHP.
The Pantheon+ data come with their errors in distance moduli in their release and the CCHP distance moduli and errors are generated as in Uddin et al. (2024) code given the entries in the light curve parameterization. We see similar distances moduli in Table 15 by our method when compared with those from Pantheon+ and with Cepheids by the CCHP. There are some differences (in 3 SNe Ia) due to the underestimate of reddening (see Figure 13), as we will discuss in the next subsection.
@@
We have modeled the errors in the distance moduli for our method adding the various contributionss. We have added in quadrature (see 4) the error in the fit of the distance modulus obtained with the MCMC, , and the covariance between the modulus and the (E(B-V)), cov (, E(B-V)), which is very low for our supernovae. We have as well the error due to a difference of stretch between the Hubble flow SNeIa and their nearby twins that goes into the distance modulus . We model it taking into account the CSP impact of the error in stretch in the distance moduli of the SNe Ia. We must say that, in our method, twins at low z and in the Hubble flow have very close stretches: within 0.02 in sBV. The light curves agree so well between low z SNe Ia and their twins in the Hubble flow (once one makes the cosmological (1+z) time broadening and magnitude dilution) that, in their plot of one over the other, they are often indistinguishable (see Appendix).
@@
To calculate the value of H0 we use the expression valid for a flat cosmology, 0:
where H0 is the Hubble constant, D is the distance to the SN Ia, c is the speed of light, z is the cosmological redshift, zHe the heliocentric redshift. The first and second order terms multiplied by z in the parenthesis involve the deceleration parameter, q0, and the cosmic jerk, j0. q0 which taking 0.3 give q0 –0.55 ( 1 - for a flat cosmology). The cosmic jerk can be taken to be j0 1 for the range of redshifts that we are using.
@@ 7.3 The value of H0
Figure 16 is the Hubble diagram of the SNe Ia in this work anchored in Cepheids (Table 15). The slope corresponds to H0 72.833 km s-1 Mpc-1. The slope of the Hubble diagram, when the SNe Ia are anchored in JAGBs (CCHP) (Table 18), is quite similar with H0 of 72.388 km s-1 Mpc-1.